OpenStudy (jetman24):

Simplify the following expression by adding and subtracting: 3√3 - 4√3 + 2√27 A. 5√3 B. 7√3 C. 10√3 D. 17√3

OpenStudy (jetman24):

medal to whoever solves this

OpenStudy (photon336):

let's simplify \[2\sqrt{27} = 2*(\sqrt{9}*\sqrt{3}) = 2*(3\sqrt{3}) = 6\sqrt{3}\] so it becomes \[3\sqrt{3}-4\sqrt{3}+6\sqrt{3}\] You would then apply order of operations addition first followed by subtraction.

OpenStudy (photon336):

@jetman24

OpenStudy (jetman24):

So what is the answer @Photon336 ?

OpenStudy (photon336):

yeah, we can't just give out direct answers. first pretend that the radical's arent there/ so we would have 3-4+6 after you figure that out put the radical three next to the result.